OpenStudy (kootra):

Evaluate the expression for x=19, y=1: x+7 ----- y+1 *PLATO sucks.. I have to TYPE EVERYTHING!!!!!!

OpenStudy (texaschic101):

(x + 7) / (y + 1) --- x = 19 and y = 1) so now we sub (19 + 7) / (1 + 1) = 26/2 which reduces to 13

OpenStudy (kootra):